OpenStudy (anonymous):

A point moves around a circle x^2 + y^2 = 117. When the point is at (6, 9), its x coordinate is increasing at the rate of 10 units per second.How fast is its y coordinate changing at that instant?

OpenStudy (anonymous):

Use implicit differentiation.

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

differentiate with respect to \(t\).

OpenStudy (anonymous):

Do you want more help?

OpenStudy (anonymous):

yes, because I keep getting wrong answer. I know it's -x/y, but I cant go further

OpenStudy (anonymous):

Both \(x\) and \(y\) are functions of \(t\).

OpenStudy (anonymous):

So what it would be?

OpenStudy (anonymous):

\[ \frac{dx^2}{dt}=\frac{dx^2}{dx}\frac{dx}{dt} = 2x\frac{dx}{dt} \]

OpenStudy (anonymous):

but what about y?

OpenStudy (anonymous):

It's completely symmetric for \(y\). So the equation is: \[ 2x\frac{dx}{dt}+2y\frac{dy}{dt}=0 \]

OpenStudy (anonymous):

but dx/dt is 10, right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and you solve for dy/dt?

OpenStudy (anonymous):

Yes.

OpenStudy (agent0smith):

And plug in (6, 9) for x and y.

OpenStudy (anonymous):

ohhhhhh,thank you guys sooooo much!